i completed 100 questions on a test. my teacher said i got 2/3 correct. how many questions did i get correct?

Found 2 solutions by jim_thompson5910, Alf101:
Answer by jim_thompson5910(35256)

You got (2/3)*100 = (2*100)/3 = 200/3 = 66.67, which rounds to 66 problems correct.

Answer by Alf101(1)

100*(2/3)=66.666 repeating which rounds up to 67% correct.